OpenStudy (mary.rojas):

Can someone walk me through this? please! Given a= i+i^3+i^5+i^7+....,which of the following represents a if the last term in the series is i^41 ? a. i b. -i c. 0 d. 1 e. -1

OpenStudy (anonymous):

You can reduce the exponents of the imaginary unit modulo 4.

OpenStudy (mary.rojas):

what do you mean "imaginary unit modulo 4."

OpenStudy (anonymous):

$$n=n+4$$ In the ring of integers modulo 4

OpenStudy (mary.rojas):

I am sorry but i was not taught that way so I dont know what your talking about

hartnn (hartnn):

ok, you know what 'i' is ?

OpenStudy (mary.rojas):

yes. but its beeen a while

OpenStudy (primeralph):



hartnn (hartnn):

ok, \(\large i= \sqrt{-1}\) so, what about i^2 =... ?

OpenStudy (mary.rojas):

-1

hartnn (hartnn):

correct! i^2 = -1 what about i^4 = (i^2)^2 =... ?

OpenStudy (mary.rojas):

1

OpenStudy (primeralph):